Please help me solve this problem!
If A=10 degrees, C=40 degrees, and side c=2, find side b.
Thank you, Natalie

Answer by Alan3354(69443)

If A=10 degrees, C=40 degrees, and side c=2, find side b.
-------------
Angle B = 180-(10+40) = 130 degrees.
Use the law of sines
sin(A)/a = sin(B)/b = sin(C)/c
side c is 2
(sin(C))/2 = (sin(B))/b
so b = 2*sin(130)/sin(40)
b = 2.3835...